OpenStudy (anonymous):

In a right traingle let L1 and L2 denote the lengths of the legs, let h denote the lengths of the hypotenuse, and let H1 and H2 denote the lengths of the segments of the hypotenuse formed by the bisector of the right angle. If L1 = 6 and L2 = 8, find the other three lengths to the nearest tenth. A) h = 10.0, H1 = 4.3 and H2 = 5.7 .. B) H = 12.2, H1 = 5.9 and H2 =6.2 .. C) H = 11.21, H1 = 4, and H2 = 5.3 .. D_ H = 9.0, H1 = 2.9, and H2 = 4.8

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

10^2 =6^2 +8^2 and 6/14 *10 =4.3 (appx)